a woman walks 250m in the direction 35 east of north. then 170m directly east. using graphical methods, find her final displacement from the starting point. compare the magnitude of her displacement with the distance she walked.

Grade:12th pass

1 Answers

Ashutosh Mohan Sharma
7 years ago
north displacement = 250 cos 30 = 216.5
east displacement = 250 sin 30 + 175 = 300
displacement = sqrt (north^2+east^2)
= 370 m
total distance = 250+175 = 425